Use the Rational Zero Theorem to find all real zeros. (Enter your answers as a comma-separated list.) 9x^3 − 17x^2 + 25x + 3 = 0
Answer #1

at states that m n-) t...tantao ithen k Zero Rational Zero Theorem :-) if the polynomial $(x) = Anxe + an- x has integer coe3 96-*) 177-*)+251-7) +3= 0 -1 81 17 81 25 q + 3 =0 = -1-17-295 + 243 81 True O N = مه zero of equation Hence lo Now using syNt. 91²-18x +27=0 Met zo n=-1 - Nowe 9x² 18x + 270 * 9(x²-2x + 3) = 6 x²-22+3=0 Hence Xa 27 (-2)=-4(1)(3) 210) cering mee datx = 2 I J 8 2. Na 2 2 2 sa d At 2 2 n = It a These are are complex root which not real. Hence, only on is real zero of 9x² -